OpenStudy (anonymous):

what is the slope between the two points (-1, -1) and (8, 0)?

OpenStudy (anonymous):

Have you see the equation: \[\huge m={y_2 - y_1 \over x_2-x_1}\]

OpenStudy (anonymous):

Where m is the slope, IE: \[\huge y=mx+b\]

OpenStudy (anonymous):

This here is the slope formula. Just plug in the coordinates. \[m = \frac{y_{1} - y_2}{x_1 - x_2}\]\[m = \frac{y_1 - y_2}{x_1 - x_2} \implies \frac{-1 - 0}{-1 - 8} \implies ?\]

OpenStudy (anonymous):

and so the answer would be 1/9 !

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Thanks!